OpenStudy (anonymous):

In triangle MNP, angle P is a right angle. If cosN = 3/5 , then sinM = 3/5 .

OpenStudy (anonymous):

sounds reasonable

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\cos(N)=\frac{3}{4}\\ \sin(M)=\frac{3}{4}\]

OpenStudy (anonymous):

sin (m) = 3/4 ??